A coil has resistance of 10 Ω and an inductance of 0.4 henry. It is connected to an AC source of 6.5 V, 30/π Hz. Find the average power consumed in the circuit.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. R = 10 Ω, L = 0.4 Henry E = 6.5 V, f = 30/π Hz Z = √R^2 + X base L^2 = √R^2 + (2πfL)^2 Power = V base rms I base rms cos ∅ = 6.5 * 6.5/Z * R/Z = 6.5 *6.5 *10/[√R^2 + (2πfL)^2]^2 = 6.5 *6.5 *10/10^2 +(2π * 30/π *0.4)^2 = 6.5 *6.5 *10/100 +576 = 0.625 = 5/8 ω